OpenStudy (safia21):

Write the standard form of the equation of the line through the given point with the given slope. through : ( 4, -2) slope = - 1

OpenStudy (anonymous):

y+2=-1(x-4)

OpenStudy (mimi_x3):

y = mx + b -2 = -1(4) + b -2 = - 4 + b b = 2 y = -x + 2

OpenStudy (anonymous):

y+2=-x-4 y=-x-4-2 y=-x-2